a. A ball is thrown horizontally at 8 m/s. (a) How fast is it moving 2 s later? In what direction is it moving?

I need an illustration and complete solution. Thank you
1
Expert's answer
2020-11-13T09:50:56-0500


a) Find the vertical velocity "v_y=g\times t = 10\times 2 = 20"  m/s.

Then the ball speed will be "v =\sqrt {v _ y ^ 2 + v _ x ^ 2} =\sqrt {20 ^ 2 + 8 ^ 2 }\approx 21.54"  m/s.

The angle of the ball trajectory in 2 seconds is at an angle to the horizon equal to "\arctg {\frac {20} {8} }\approx 68\degree" .



b) Since the distance travelled (where "1524"  meters "\approx 5000"  feet) is "S =\frac {g\times t ^ 2} {2}" , therefore "t =\sqrt {\frac {2\times S} {g}} =\sqrt {\frac {2\times 1524} {10} }\approx17.46"  s

Then the vertical velocity is "v_y=10\times 17.46 = 174.6"  m/s

"300"  miles/hours "\approx134.08"  m/s.

"v_{total}=\sqrt{174.6^2+134.08^2}\approx 220.14"  m/s